If you put $4,000 in a savings account that pays interest at the rate of 4%, compounded annually, how much will you have in 5 years? (Hint, use the future value formula.) How much interest will you earn during the 5 years? If you put %4,000 each year into a savings account that pays interest at the rate of 4% a year, how much would you have after 5 years?
Answer by rfer(16322)

FV=4000(1.04)^5
FV=4000(1.21665)
FV=$4866.60
-------------
4866.60-4000=$866.69 interest